OpenStudy (anonymous):

using integration by substitution what is the integral of 2xe^(x^2-4)

OpenStudy (lalaly):

let u=x^2-4 du=2xdx so the integral becomes\[\int\limits{e^udu}\]

OpenStudy (anonymous):

thxxx

OpenStudy (lalaly):

:):)